The shareholders' equity of Nick Co. includes the items shown below. The board of directors of Nick declared cash dividends of $4.5 million, $10.5 million, and $52.5 million in each of its first 3 years of operation: 2016, 2017, and 2018, respectively.

Common stock, $1 par, 50,000,000 shares outstanding
Preferred stock, 6%, $100 par, 1,000,000 shares outstanding

Required:
Determine the amount of dividends per share on preferred and common stock for each of the three years. The preferred stock is noncumulative and nonparticipating. (Round final answers to 2 decimal places.)
Answer #1
Preferred Common
2016 4500000 0
2017 7500000 3000000
2018 6000000 46500000

workings

Year 2016
Preferred Stock dividend =1000000*100*6% 6000000
Dividend Paid to Preferred stock dividend 4500000
Preferred stock dividend arrear (600000-450000) 1500000
Year 2017
Preferred Stock dividend =1000000*100*6% 6000000
Preferred stock dividend arrear (600000-450000) 1500000
Dividend Paid to Preferred stock dividend 7500000
Total dividend declared 10500000
Dividend Available to common stock holder 1050000-7500000 3000000
Year 2018
Preferred Stock dividend =1000000*100*6% 6000000
Total dividend declared 52500000
Dividend Available to common stock holder 52500000-6000000 46500000